# UniFlow: Towards Zero-Shot LiDAR Scene Flow for Autonomous Vehicles via Cross-Domain Generalization

### v1

@@ Abstract
LiDAR scene flow is the task of estimating per-point 3D motion between consecutive point clouds. Recent methods achieve centimeter-level accuracy on popular autonomous vehicle (AV) datasets, but are typically only trained and evaluated on a single sensor. In this paper, we aim to learn general motion priors that transfer to diverse and unseen LiDAR sensors. However, prior work in LiDAR semantic segmentation and 3D object detection demonstrate that naively training on multiple datasets yields worse performance than single dataset models. Interestingly, we find that this conventional wisdom does not hold for motion estimation, and that state-of-the-art scene flow methods greatly benefit from cross-dataset training. We posit that low-level tasks such as motion estimation may be less sensitive to sensor configuration; indeed, our analysis shows that models trained on fast-moving objects (e.g., from highway datasets) perform well on fast-moving objects, even across different datasets. Informed by our analysis, we propose UniFlow, a family of feedforward models that unifies and trains on multiple large-scale LiDAR scene flow datasets with diverse sensor placements and point cloud densities. Our frustratingly simple solution establishes a new state-of-the-art on Waymo and nuScenes, improving over prior work by 5.1% and 35.2% respectively. Moreover, UniFlow achieves state-of-the-art accuracy on unseen datasets like TruckScenes, outperforming prior TruckScenes-specific models by 30.1%. We include additional visuals on our project page.
RGB |
Sensor |
BEV |
@@ 1 Introduction
Recent LiDAR scene flow methods achieve remarkable performance on popular autonomous vehicle (AV) datasets [vedder2024neural, zeroflow, zhang2025deltaflow, khatri2024can] like Argoverse 2, nuScenes, and Waymo [sun2020scalability, nuscenes, Argoverse2_2021]. Notably, current state-of-the-art approaches train and evaluate on each dataset independently and do not attempt to generalize to unseen sensor configurations. Intuitively, since each dataset has different sensor placements and point cloud densities, training separate models for each domain seems reasonable (Figure 1). For example, Argoverse 2 (AV2) [Argoverse2_2021] uses two out-of-phase 32 beam LiDARs, nuScenes [nuscenes] uses one 32 beam sensor, and Waymo [sun2020scalability] uses a custom 64 beam sensor. Indeed, it is notoriously challenging to train multi-dataset models for LiDAR-based 3D object detection [wang2020train, soum2023mdt3d] and semantic segmentation [kim2024rethinking, saltori2023walking] that improve over single dataset models. Interestingly, our analysis challenges this conventional wisdom for LiDAR-based scene flow. Unlike “high-level” tasks like LiDAR semantic segmentation or 3D object detection that require human-defined taxonomies [lambert2020mseg], we argue that LiDAR scene flow models capture universal “low-level” geometric features that should, in principle, generalize better across sensors.
@@
Cross-Domain Generalization for Low-Level Vision. Notably, we observe similar trends in RGB-based low-level vision tasks, where models learn generalizable geometric representations that transfer well across datasets. For example, FlowNet [dosovitskiy2015flownet, ilg2017flownet] and RAFT [teed2020raft] show that optical flow models trained on synthetic datasets (e.g. FlyingChairs, FlyingThings3D, Sintel) generalize surprisingly well to casually captured videos. More recently, zero-shot monocular depth estimators like MoGe [wang2025moge] and monocular scene flow estimators like ZeroMSF [liang2025zero] demonstrate that training on diverse datasets significantly improves zero-shot performance. Therefore, we posit that low-level LiDAR-based scene flow may similarly generalize to different sensors if trained at scale.
@@
To better understand the problem, we first train dataset-specific models for AV2 [Argoverse2_2021], Waymo [sun2020scalability] and nuScenes [nuscenes] and evaluate zero-shot cross-domain performance. Surprisingly, Flow4D [kim2024flow4d] trained on Waymo achieves similar Dynamic Mean EPE on AV2 as Flow4D trained on AV2. Importantly, Waymo’s 64 beam LiDAR has considerably more point density than AV2’s two out-of-phase 32-beam LiDARs (Fig. 1), indicating that point density does not significantly impact LiDAR scene flow generalization. Further, Flow4D (Waymo) achieves lower EPE (i.e. better performance) than Flow4D (AV2) on fast moving objects in AV2 since there are more fast movers in Waymo than in AV2 (Table 1). This suggests that the accuracy of motion prediction is highly correlated with object velocity (Figure 2); in order to accurately predict fast motion, one should train on fast motion. This suggests that the key to generalization is training on a diverse combination of datasets with a wide range of object velocities.
@@
Towards Zero-Shot LiDAR Scene Flow. We argue that scaling up LiDAR-based scene flow methods will be a key enabler for 3D motion understanding and dynamic reconstruction in diverse environments. To this end, we retrain state-of-the-art methods on supervised data from AV2, Waymo, and nuScenes and demonstrate significant improvements for both in-distribution and out-of-distribution generalization. We denote models trained with multiple datasets as UniFlow. To the best of our knowledge, UniFlow is the first to demonstrate the benefits of cross-domain training for LiDAR scene flow. Further, UniFlow models demonstrate remarkable zero-shot accuracy on TruckScenes [fent2024man], outperforming prior dataset-specific models by 30.1% (Table 6). Lastly, we find that cross-domain training on near-range (i.e 35 meters from the ego-vehicle) points improves scene flow generalization to unseen far-range points (Table 7). Intuitively, sparse LiDAR points near the ego-vehicle mimic the distribution of dense LiDAR points far away from the vehicle [peri2023empirical].
Why Does Multi-Dataset Training Work? Our experiments suggest that multi-dataset training improves LiDAR scene flow performance because scene flow is inherently class-agnostic. This allows us to avoid complications arising from dataset-specific label definitions (Table 8). For example, nuScenes defines bicycle as excluding the rider, while Waymo includes the rider [madan2023revisiting, robicheaux2025roboflow100]. Such label ambiguity poses a significant challenge for semantic tasks, potentially explaining why prior LiDAR-based semantic tasks like 3D object detection and semantic segmentation do not benefit from training on multiple datasets. Furthermore, we find that with careful augmentations, we can simulate diverse sensor configurations that improve cross-dataset generalization (Figure 4). For example, dropping out LiDAR beams from a Waymo point cloud makes the data look more like a point cloud from AV2 or nuScenes, while augmenting the height of a point cloud makes a sensor mounted on a sedan look like it was mounted on a truck.
Contributions. We present three major contributions. First, we highlight that dataset-specific scene flow models already show signs of cross-domain generalization, motivating our investigation into multi-dataset training. Next, we demonstrate that re-training existing scene flow methods on diverse, large-scale data yields state-of-the-art performance on nuScenes and Waymo, improving over our baselines by 35.2% and 5.1% respectively. Finally, we show that UniFlow outperforms dataset-specific models by 30.1% on TruckScenes, challenging conventional wisdom on LiDAR-based cross-domain generalization.
2 Related Work
@@
Cross-Domain Generalization for LiDAR is a long-standing challenge in LiDAR-based perception. Prior works in 3D object detection [malic2025gblobs, hegde2025multimodal] and semantic segmentation [kim2024rethinking, caunes2025seg_3d_by_pc2d, xu2025superflowpp] show that models trained on multiple datasets often perform worse than dataset-specific models. This performance drop can be largely attributed to diverse sensor hardware (e.g., number of beams, scan pattern, point cloud density), environmental conditions (e.g., weather, geography), and sensor placement across datasets. To address these challenges, several methods have introduced data augmentation strategies [sun2024empirical], such as random point dropping [wang2021pointaugmenting] and object scaling [cen2022open], while others [Liu_2024_CVPR, michele2024saluda] have proposed unsupervised domain adaptation techniques that align feature distributions across datasets. However, such data augmentation strategies primarily address the geometric domain gap between datasets. For motion-centric tasks such as scene flow, we argue that understanding the distribution of object velocities between datasets represents a critical yet overlooked axis for addressing cross-domain generalization. To this end, our work presents the first systematic study of this velocity domain gap.
@@ 3 UniFlow: Towards Zero-Shot LiDAR Scene Flow for Autonomous Vehicles
UniFlow aims to learn general motion priors that transfer to diverse and unseen LiDAR sensors by training a single scene flow model across diverse autonomous driving datasets. In this section, we describe the process of unifying four popular AV datasets and our training strategy.
Dataset Unification. We unify four widely used AV datasets to facilitate multi-dataset training. Importantly, we standardize LiDAR sensor frame rates and annotation frequencies to ensure that the time interval between LiDAR sweeps is consistent across datasets. This step is crucial because state-of-the-art scene flow methods parameterize motion as the displacement between two consecutive LiDAR frames. Specifically, we down-sample high-frequency datasets like nuScenes (which provides LiDAR sweeps at 20 Hz) to 10 Hz to match AV2 and Waymo. For datasets with sparse scene flow annotations (e.g., nuScenes annotates tracks at 2 Hz), we train solely on the annotated frames and their real adjacent scans. This ensures that we do not introduce any label noise or motion artifacts due to interpolation. Lastly, we apply LineFit [himmelsbach2010fast] for ground point removal for all datasets.
Unified Multi-Dataset Training. We retrain existing state-of-the-art scene flow architectures (SSF [khoche2025ssf], Flow4D [kim2024flow4d], and Flow [zhang2025deltaflow]) on the unified dataset mixture of nuScenes, AV2, and Waymo. Importantly, we do not re-weight dataset frequencies or apply dataset-specific augmentations. Instead, we employ a minimal yet effective augmentation set including height randomization to account for sensor placement and LiDAR ray dropping to simulate varying LiDAR point densities. This unified training paradigm encourages models to learn domain-agnostic geometric priors rather than memorizing dataset-specific motion patterns, resulting in significant zero-shot generalization improvements.
@@
| Val | 75 | 14,625 / 2,925 |
Overall Performance |
Rare Vehicles |
Long Range ( m) |
(a) Flow
(b) Flow (UniFlow)
(c) Ground truth
@@ 4 Experiments
Datasets. We train and evaluate UniFlow on four large-scale autonomous driving datasets to rigorously evaluate cross-domain generalization to diverse sensor configurations, ego-vehicle platforms, and driving scenarios (Table 2). AV2 and nuScenes collect data in dense urban environments with 32-beam LiDARs. Waymo uses a custom LiDAR sensor to collect denser point clouds from urban and suburban environments with more varied driving dynamics. We evaluate zero-shot generalization using TruckScenes [fent2024man] since it collects data in different environments (high-speed highway scenes vs. urban and suburban scenes), with different sensors (32-beam LiDARs vs. 64-beam LiDARs), using a different vehicle type (truck vs. sedan). These four datasets allow us to systematically analyze the impact of point density (32-beam vs. 64-beam), ego-vehicle perspective (car vs. truck), and velocity (urban vs. highway) on cross-domain generalization.
Baselines. We compare UniFlow against both self-supervised and fully-supervised methods: NSFP [li2021neural], FastNSF [li2023fast], ICPFlow [lin2024icp], SeFlow [zhang2024seflow], SeFlow++ [zhang2025himo], EulerFlow [vedder2024neural], VoteFlow [lin2025voteflow], TrackFlow [khatri2024can], DeFlow [zhang2024deflow], SSF [khoche2025ssf], Flow4D [kim2024flow4d] and Flow [zhang2025deltaflow]. To ensure fairness, Argoverse 2 test-set results are obtained directly from the public leaderboard. We use Waymo, nuScenes and TruckScenes baselines as implemented in OpenSceneFlow111https://github.com/KTH-RPL/OpenSceneFlow with standard training hyperparameters.
@@
| Flow4D (UniFlow, Ours) | 2.07 | 4.50 | 1.31 | 0.40 | 0.132 | 0.072 | 0.131 | 0.218 | 0.107 |
@@
| Flow4D (UniFlow, Ours) | 1.60 | 3.83 | 0.68 | 0.30 | 0.191 | 0.064 | 0.400 | 0.110 |
@@
| Flow4D (UniFlow, Ours) | 3.01 | 7.28 | 1.47 | 0.28 | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 |
@@
| Flow4D (UniFlow, Ours) | 23.59 | 68.41 | 1.78 | 0.57 | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 |
| Flow (UniFlow, Ours) | 15.76 | 45.76 | 1.02 | 1.32 | 0.283 | 0.101 | 0.293 | 0.513 | 0.226 |
@@
Comparison to State-of-the-Art Methods. We demonstrate the effectiveness of the UniFlow framework by retraining SSF, Flow4D, and Flow on multiple datasets and compare their performance against recent approaches on popular scene flow benchmarks (Tables 3, 4, and 5). On the Argoverse 2 test set, Flow4D (UniFlow) improves over Flow4D by 8.97% and SSF (UniFlow) improves over SSF by 13.81% in Dynamic Mean EPE. Moreover, Flow4D (UniFlow) achieves near-parity with EulerFlow at a fraction of the runtime. We see similar improvements between Flow4D and SSF and their UniFlow counterparts on the Waymo (Table 4) and nuScenes (Table 5) validation sets. Interestingly, SSF (UniFlow) outperforms Flow4D (UniFlow) on nuScenes, suggesting that the SSF architecture is naturally better suited for processing sparse point clouds.
@@
TruckScenes Zero-Shot Performance. We demonstrate UniFlow’s strong zero-shot generalization capabilities with extensive evaluations on TruckScenes (Table 6). All baseline methods (except UniFlow) were trained on TruckScenes. Despite never being explicitly trained on high speed highway scenes, Flow4D (UniFlow) significantly outperforms the dataset-specific Flow4D on Dynamic Mean EPE. Notably, Flow4D (UniFlow) was only trained on urban driving scenes collected from a sedan, indicating that our approach can successfully generalize across significant domain gaps. Table 6 also highlight the importance of the speed normalized and bucketed EPE metric [khatri2024can], as Threeway-EPE results don’t adequately measure Flow4D’s (UniFlow) significant improvements on pedestrian and VRU performance.
@@
Analysis of Long-Range Generalization. To evaluate whether unified training produces range-robust representations, we evaluate how models trained on short-range LiDAR data transfer to longer-range settings without retraining. Specifically, we extend the voxelization range during inference while keeping the voxel size identical to the original trained model. This change alters only the spatial coverage of the dynamic voxelization module, requiring a simple weight transfer step to adapt to the new grid dimensions while preserving all learned features. As shown in Table 7, UniFlowLR consistently outperforms FlowLR across all distance intervals on AV2, reducing Dynamic Mean EPE by 12.7% on average and up to 20.0% in the 75–100 m range. This indicates that unified multi-dataset training produces representations that remain robust even when the perception range is significantly expanded. On TruckScenes, UniFlowLR,ZS achieves a remarkable 40.6% reduction in overall error compared to FlowLR, with improvements exceeding 50% in the 50–100 m range, despite never being trained on this dataset. These results highlight that the representations learned through unified multi-dataset training naturally scale to larger spatial extents.
4.1 Ablation Studies
Analysis of Semantic Cross-Domain Generalization. We evaluate the benefits unified training on both geometric and semantic understanding by extending Flow4D with a multi-task semantic head. This new Flow4D variant simultaneously predicts scene flow and per-point semantics using the same backbone features. Since all datasets follow different taxonomies, we map their annotations to four coarse categories (e.g. car, other, pedestrian, and VRU) and jointly train the model with an additional class-balanced cross-entropy loss. As shown in Table 8, adding the semantic head has almost no effect on scene flow accuracy, indicating that the auxiliary task does not interfere with motion estimation. Consistent with our original findings, unified multi-dataset training continues to improve both in-domain and zero-shot performance for scene flow. However, training the semantic prediction head on the unified dataset yields mixed results. Models trained on unified datasets show small in-domain gains, but perform substantially worse than dataset-specific model when evaluated on unseen domains. This supports our hypothesis that dataset unification primarily benefits low-level geometric representations, while high-level semantics remain sensitive to dataset-specific biases.
@@ 4.1 Ablation Studies
| Train | Test | 3-Way (cm) | Dyn. Norm. | mIoU | Accuracy | CAR | OTHER | PED | VRU |
@@
Analysis of Training Recipe. We ablate the impact of multi-dataset training, data augmentation, and model scaling on zero-shot TruckScenes performance in Table 9. All baselines are trained on Waymo only to ensure zero-shot evaluation. All model architectures show clear gains from multi-dataset training. Notably, despite training on a dataset three times larger, data augmentation continues to show substantial improvements, indicating that further scaling may yield additional performance gains.

### v2

# UniFlow: Zero-Shot LiDAR Scene Flow for Autonomous Driving

@@ Abstract
LiDAR scene flow is the task of estimating per-point 3D motion between consecutive point clouds. Recent methods achieve centimeter-level accuracy on popular autonomous vehicle (AV) datasets, but are typically only trained and evaluated on a single sensor. In this paper, we aim to learn general motion priors that transfer to diverse and unseen LiDAR sensors. However, prior work in LiDAR semantic segmentation and 3D object detection demonstrate that naively training on multiple datasets yields worse performance than single dataset models. Interestingly, we find that this conventional wisdom does not hold for motion estimation, and that state-of-the-art scene flow methods greatly benefit from cross-dataset training without architectural modification. We posit that low-level tasks such as motion estimation may be less sensitive to sensor configuration; indeed, our analysis shows that models trained on fast-moving objects (e.g., from highway datasets) perform well on fast-moving objects, even across different datasets. Informed by our analysis, we propose UniFlow, a feedforward model that unifies and trains on multiple large-scale LiDAR scene flow datasets with diverse sensor placements and point cloud densities. Our frustratingly simple solution establishes a new state-of-the-art on Waymo and nuScenes, improving over prior work by 5.1% and 35.2% respectively. Moreover, UniFlow achieves state-of-the-art accuracy on unseen datasets like TruckScenes and AEVAScenes, outperforming prior dataset-specific models by 30.1% and 22.5% respectively. See our project page for additional visuals.
RGB
Sensor
BEV
Argoverse 2
Waymo
nuScenes
TruckScenes
@@ 1 Introduction
Recent LiDAR scene flow methods achieve remarkable performance on popular autonomous vehicle (AV) datasets [vedder2024neural, zeroflow, zhang2025deltaflow, khatri2024can] like Argoverse 2, nuScenes, and Waymo [sun2020scalability, nuscenes, Argoverse2_2021]. Notably, current state-of-the-art approaches train and evaluate on each dataset independently and do not attempt to generalize to unseen sensor configurations. Intuitively, since each dataset has different sensor placements and point cloud densities, training separate models for each domain seems reasonable (Figure 1). For example, Argoverse 2 (AV2) [Argoverse2_2021] uses two out-of-phase 32 beam LiDARs, nuScenes [nuscenes] uses one 32 beam sensor, and Waymo [sun2020scalability] uses a custom 64 beam sensor. Indeed, it is notoriously challenging to train multi-dataset models for LiDAR-based 3D object detection [wang2020train, soum2023mdt3d] and semantic segmentation [kim2024rethinking, saltori2023walking, liu2024multi] that improve over single dataset models. Notably, prior works that do train on multiple datasets introduce bespoke architectures to explicitly address the problem of cross-domain alignment. Our key contribution is an empirical analysis that shows this does not hold for LiDAR-based scene flow; we simply expose standard architectures without modification to different sensor data and see a consistent improvement. We posit that “high-level” tasks like LiDAR semantic segmentation or 3D object detection are challenging to train across domains because they rely on human-defined taxonomies [lambert2020mseg] that may be inconsistent across datasets [madan2023revisiting]. Instead, scene flow models appear to capture properties of the underlying physical 4D world that hold regardless of LiDAR sensors observing it.
@@
Cross-Domain Generalization for Low-Level Vision. We observe similar trends in RGB-based low-level vision tasks, where models learn generalizable geometric representations that transfer well across datasets. For example, FlowNet [dosovitskiy2015flownet, ilg2017flownet] and RAFT [teed2020raft] show that optical flow models trained on synthetic datasets (e.g. FlyingChairs, FlyingThings3D, Sintel) generalize surprisingly well to casually captured videos. More recently, zero-shot monocular depth estimators like MoGe [wang2025moge] and monocular scene flow estimators like ZeroMSF [liang2025zero] demonstrate that training on diverse datasets significantly improves zero-shot performance. Therefore, we posit that low-level LiDAR-based scene flow may similarly generalize to different sensors if trained at scale.
@@
To better understand this problem, we first train dataset-specific models for AV2 [Argoverse2_2021], Waymo [sun2020scalability] and nuScenes [nuscenes] and evaluate zero-shot cross-domain performance. Surprisingly, Flow4D [kim2024flow4d] trained on Waymo achieves similar Dynamic Mean EPE on AV2 as Flow4D trained on AV2. Importantly, Waymo’s 64 beam LiDAR is considerably more dense than AV2’s two out-of-phase 32-beam LiDARs (Figure 1), indicating that point density does not significantly impact LiDAR scene flow generalization. Further, Flow4D (Waymo) achieves lower EPE (i.e. better performance) than Flow4D (AV2) on fast moving objects in AV2 since there are more fast movers in Waymo than in AV2 (Table 1). This suggests that the accuracy of motion prediction is highly correlated with object velocity (Figure 2); in order to accurately predict fast motion, one should train on fast motion. This suggests that the key to LiDAR scene flow generalization is training on a diverse combination of datasets with a wide range of object velocities.
@@
Towards Zero-Shot LiDAR Scene Flow. We argue that scaling up LiDAR-based scene flow methods will be a key enabler for 3D motion understanding and dynamic reconstruction in diverse environments. To this end, we retrain state-of-the-art methods without modification on supervised data from AV2, Waymo, and nuScenes and demonstrate significant improvements for both in-distribution and out-of-distribution generalization. We denote models trained with multiple datasets as UniFlow. To the best of our knowledge, UniFlow is the first to demonstrate the benefits of cross-domain training for LiDAR scene flow. Further, UniFlow models demonstrate remarkable zero-shot accuracy on TruckScenes [fent2024man] (an autonomous truck dataset collected at highway speeds) and AEVAScenes (a small-scale dataset with a novel FMCW LiDAR sensor), outperforming prior dataset-specific models by 30.1% (Table 6) and 22.5% (Table 7), respectively. Lastly, we find that cross-domain training on near-range (i.e 35 meters from the ego-vehicle) points improves scene flow generalization to unseen far-range points (Table 8). We hypothesize that sparse LiDAR points near the ego-vehicle mimic the distribution of dense LiDAR points far away from the vehicle [peri2023empirical].
Why Does Multi-Dataset Training Work? Our experiments suggest that multi-dataset training improves LiDAR scene flow performance because scene flow is inherently class-agnostic. This allows us to avoid complications arising from dataset-specific label definitions (Table 9). For example, nuScenes defines bicycle as excluding the rider, while Waymo includes the rider [madan2023revisiting, robicheaux2025roboflow100]. Such label ambiguity poses a significant challenge for semantic tasks, potentially explaining why naively combining datasets for LiDAR-based 3D object detection and semantic segmentation yields worse performance than single-dataset models. Furthermore, we find that with careful augmentations, we can simulate diverse sensor configurations that improve cross-dataset generalization (Figure 4). For example, dropping out LiDAR beams from a Waymo point cloud makes the data look more like a nuScenes point cloud, while augmenting point cloud height makes a sensor mounted on a sedan look like it was mounted on a truck.
Contributions. We present three major contributions. First, we highlight that dataset-specific scene flow models already show signs of cross-domain generalization, motivating our investigation into multi-dataset training. Next, we demonstrate that re-training existing scene flow methods on diverse, large-scale data yields state-of-the-art performance on nuScenes and Waymo, improving over our baselines by 35.2% and 5.1% respectively. Finally, we show that UniFlow generalizes well to unseen sensors, outperforming dataset-specific models by 30.1% on TruckScenes, and by 22.5% on AEVAScenes.
Overall
Performance
Rare
Vehicles
Long Range
( m)
(a) Flow
(b) Flow (UniFlow)
(c) Ground truth
2 Related Works
@@
Cross-Domain Generalization for LiDAR is a long-standing challenge in LiDAR-based perception. Prior works in 3D object detection [malic2025gblobs, hegde2025multimodal] and semantic segmentation [kim2024rethinking, caunes2025seg_3d_by_pc2d, xu2025superflowpp] show that naively training models on multiple LiDAR datasets often performs worse than models trained on a single dataset. This performance drop can be largely attributed to diverse sensor hardware (e.g., number of beams, scan pattern, point cloud density), environmental conditions (e.g., weather, geography), and sensor placement across datasets. Often, such methods also design bespoke architectures to address cross-domain generalization. In addition, several methods have introduced data augmentation strategies [sun2024empirical], such as random point dropping [wang2021pointaugmenting] and object scaling [cen2022open], while others [Liu_2024_CVPR, michele2024saluda] have proposed unsupervised domain adaptation techniques that align feature distributions across datasets. However, such data augmentation strategies primarily address the geometric domain gap between datasets. For motion-centric tasks such as scene flow, we argue that understanding the distribution of object velocities between datasets represents a critical yet overlooked axis for addressing cross-domain generalization. To this end, our work presents the first systematic study of this velocity domain gap.
@@ 3 UniFlow: Towards Zero-Shot LiDAR Scene Flow for Autonomous Vehicles
UniFlow aims to learn general motion priors that transfer to diverse and unseen LiDAR sensors by training a single scene flow model across diverse autonomous driving datasets. In this section, we describe the process of unifying five popular AV datasets and our training strategy.
Dataset Unification. We unify five widely used AV datasets to facilitate multi-dataset training. Importantly, we standardize LiDAR sensor frame rates and annotation frequencies to ensure that the time interval between LiDAR sweeps is consistent across datasets. This step is crucial because state-of-the-art scene flow methods parameterize motion as the displacement between two consecutive LiDAR frames. Specifically, we down-sample high-frequency datasets like nuScenes (which provides LiDAR sweeps at 20 Hz) to 10 Hz to match AV2 and Waymo. For datasets with sparse scene flow annotations (e.g., nuScenes annotates tracks at 2 Hz), we train solely on the annotated frames and their real adjacent scans. This ensures that we do not introduce any label noise or motion artifacts due to interpolation. Lastly, we apply LineFit [himmelsbach2010fast] for ground point removal for all datasets.
Unified Multi-Dataset Training. We retrain existing state-of-the-art scene flow architectures (SSF [khoche2025ssf], Flow4D [kim2024flow4d], and Flow [zhang2025deltaflow]) on the unified dataset mixture of nuScenes, AV2, and Waymo. Importantly, we do not re-weight dataset frequencies or apply dataset-specific augmentations. Instead, we employ a minimal yet effective augmentation set including height randomization to account for sensor placement (, ) and LiDAR ray dropping to simulate varying LiDAR point densities (alternating odd/even rays, ). Ray dropping is not performed on nuScenes due to its inherent sparsity. Further, we do not discard any scans and use dataset provided ego-motion for motion compensation. We do not temporally align sensor measurements between datasets beyond frame rate. This unified training paradigm encourages models to learn domain-agnostic geometric priors rather than memorizing dataset-specific motion patterns, resulting in significant zero-shot generalization improvements. See Appendix 0.A for more details.
@@
| Val | 75 | 14,625 / 2,925 | ||||
| AEVAScenes | Val | 65 | 6,500 | 6 FMCW | Car | Urban / Highway |
@@ 4 Experiments
In this section, we evaluate UniFlow on Argoverse 2, Waymo, nuScenes, TruckScenes, and AEVAScenes. We ablate the impact of dataset characteristics on model performance, and identify scaling trends. We encourage the reader to see the supplement for additional analysis and visuals.
@@
| Flow4D-XL (UniFlow, Ours) | 2.07 | 4.50 | 1.31 | 0.40 | 0.132 | 0.072 | 0.131 | 0.218 | 0.107 |
@@
| Flow4D-XL (UniFlow, Ours) | 1.60 | 3.83 | 0.68 | 0.30 | 0.191 | 0.064 | 0.400 | 0.110 |
@@
| Flow4D-XL (UniFlow, Ours) | 3.01 | 7.28 | 1.47 | 0.28 | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 |
@@
| Flow4D-XL (UniFlow, Ours) | 23.59 | 68.41 | 1.78 | 0.57 | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 |
| Flow (UniFlow, Ours) | 16.04 | 45.76 | 1.02 | 1.32 | 0.283 | 0.101 | 0.293 | 0.513 | 0.226 |
| Method | Three-way EPE (cm) | Dynamic Bucket-Norm. | |||||||
| Mean | FD | FS | BS | Mean | Car | Other | Pedestr. | VRU | |
| SSF (AV2) | 32.13 | 92.92 | 2.09 | 1.39 | 0.822 | 0.822 | 0.851 | 0.858 | 0.859 |
| SSF (Waymo) | 20.59 | 59.38 | 1.86 | 0.53 | 0.811 | 0.513 | 0.957 | 0.877 | 0.899 |
| SSF (nuScenes) | 31.04 | 88.68 | 3.08 | 1.36 | 0.772 | 0.689 | 0.865 | 0.775 | 0.759 |
| SSF (UniFlow, Ours) | 10.80 | 29.76 | 2.05 | 0.60 | 0.544 | 0.239 | 0.686 | 0.612 | 0.639 |
| Flow4D (AV2) | 16.13 | 43.20 | 3.19 | 2.01 | 0.419 | 0.317 | 0.434 | 0.491 | 0.433 |
| Flow4D (Waymo) | 8.14 | 18.89 | 4.03 | 1.50 | 0.408 | 0.152 | 0.560 | 0.491 | 0.560 |
| Flow4D (nuScenes) | 25.61 | 70.06 | 3.86 | 2.93 | 0.535 | 0.512 | 0.678 | 0.459 | 0.489 |
| Flow4D-XL (UniFlow, Ours) | 6.14 | 13.16 | 3.89 | 1.37 | 0.338 | 0.114 | 0.312 | 0.480 | 0.448 |
| Flow (AV2) | 15.63 | 43.37 | 2.73 | 0.79 | 0.444 | 0.302 | 0.560 | 0.460 | 0.453 |
| Flow (Waymo) | 7.01 | 16.27 | 3.59 | 1.19 | 0.391 | 0.123 | 0.648 | 0.426 | 0.426 |
| Flow (nuScenes) | 21.80 | 49.91 | 4.12 | 1.56 | 0.457 | 0.396 | 0.596 | 0.435 | 0.402 |
| Flow (UniFlow, Ours) | 6.65 | 13.17 | 5.16 | 1.62 | 0.303 | 0.112 | 0.359 | 0.398 | 0.344 |
Datasets. We train and evaluate UniFlow on five large-scale autonomous driving datasets to rigorously evaluate cross-domain generalization to diverse sensor configurations, ego-vehicle platforms, and driving scenarios (Table 2). AV2 and nuScenes collect data in dense urban environments with 32-beam LiDARs. Waymo uses a custom LiDAR sensor to collect denser point clouds from urban and suburban environments with more varied driving dynamics. We evaluate zero-shot generalization using TruckScenes [fent2024man] since it collects data in different environments (high-speed highway scenes vs. urban and suburban scenes), with different sensors (32-beam LiDARs vs. 64-beam LiDARs), using a different vehicle type (truck vs. sedan). We also evaluate zero-shot generalization using AEVAScenes since it uses a novel FMCW LiDAR. These five datasets allow us to systematically analyze the impact of point density (32-beam vs. 64-beam), ego-vehicle perspective (car vs. truck), and velocity (urban vs. highway) on cross-domain generalization.
Baselines. We compare UniFlow against both self-supervised and fully-supervised methods: NSFP [li2021neural], FastNSF [li2023fast], ICPFlow [lin2024icp], SeFlow [zhang2024seflow], SeFlow++ [zhang2025himo], EulerFlow [vedder2024neural], VoteFlow [lin2025voteflow], TrackFlow [khatri2024can], DeFlow [zhang2024deflow], SSF [khoche2025ssf], Flow4D [kim2024flow4d] and Flow [zhang2025deltaflow]. We design a new model called Flow4D-XL, which simply adds more parameters to Flow4D to address model underfitting and improve its scaling behavior (Appendix 0.A). To ensure fairness, Argoverse 2 test-set results are obtained directly from the public leaderboard. We use Waymo, nuScenes and TruckScenes baselines as implemented in OpenSceneFlow111https://github.com/KTH-RPL/OpenSceneFlow with standard training hyperparameters. We include additional implementation details in the supplement.
@@
Comparison to State-of-the-Art Methods. We demonstrate the effectiveness of the UniFlow framework by retraining SSF, Flow4D, and Flow on multiple datasets and compare their performance against recent approaches on popular scene flow benchmarks (Tables 3, 4, and 5). On the Argoverse 2 test set, Flow4D-XL (UniFlow) improves over Flow4D by 9.0% and SSF (UniFlow) improves over SSF by 13.8% in Dynamic Mean EPE. Moreover, Flow4D-XL (UniFlow) achieves near-parity with EulerFlow at a fraction of the runtime. We see similar improvements between Flow4D and SSF and their UniFlow counterparts on the Waymo (Table˜4) and nuScenes (Table˜5) validation sets. Interestingly, SSF (UniFlow) outperforms Flow4D-XL (UniFlow) on nuScenes, suggesting that the SSF architecture is naturally better suited for processing sparse point clouds.
@@
TruckScenes Zero-Shot Performance. We demonstrate UniFlow’s strong zero-shot generalization capabilities with extensive evaluations on TruckScenes (Table 6). All baseline methods (except UniFlow) were trained on TruckScenes. Despite never being explicitly trained on high speed highway scenes, Flow4D-XL (UniFlow) significantly outperforms the dataset-specific Flow4D on Dynamic Mean EPE. Notably, Flow4D-XL (UniFlow) was only trained on urban driving scenes collected from a sedan, indicating that our approach can successfully generalize across significant domain gaps. Table 6 also highlight the importance of the speed normalized and bucketed EPE metric [khatri2024can], as three-way EPE results don’t adequately measure Flow4D-XL’s (UniFlow) significant improvements on pedestrian and VRU performance. We ablate this further in Appendix 0.E.
@@
| Train | Test | Three-Way (cm) | Dyn. Norm. | mIoU | Accuracy | CAR | OTHER | PED | VRU |